f(x)=x2-7x+10;
how do I find
a.) f(0)
b.) f(5)
c.) f(1)
I really don't understand, could you please break it down in laymans terms for me. I need help badly.
Nina

Found 2 solutions by AnlytcPhil, tutorcecilia:
Answer by AnlytcPhil(1116)

f(x) = x² - 7x + 10; 

how do I find

a.) f(0)
b.) f(5)
c.) f(1)  

I really don't understand, could you please break it down in 
laymans terms for me. I need help badly. 

Nina

This is probably one of the easiest thing you'll ever learn in 
algebra.  As you'll see, it's just "plug in for x and simplify":

(a)

To find f(0), just substitute 0 everywhere x is and then 
simplify the
right side:

f(x) = x² - 7x + 10

Now substitute a 0 everywhere there is an x:

f(0) = 0² - 7(0) + 10

Now simplify the right side:

f(0) = 0 - 0 + 10

f(0) = 10

----------------------------------

(b)

To find f(5), just substitute 5 everywhere x is and then 
simplify the right side:

f(x) = x² - 7x + 10

Now substitute a 5 everywhere there is an x:

f(5) = 5² - 7(5) + 10

Now simplify the right side:

f(5) = 25 - 35 + 10

f(5) = 0

----------------------------------

(c)

To find f(1), just substitute 1 everywhere x is and then 
simplify the right side:

f(x) = x² - 7x + 10

Now substitute a 1 everywhere there is an x:

f(1) = 1² - 7(1) + 10

Now simplify the right side:

f(1) = 1 - 7 + 10

f(1) = 4

See?  It's just "plug-in and simplify". What could be easier?

Edwin

Answer by tutorcecilia(2144)

Start with the basic equation "f(x)=x^2-7x+10" which is the frame-work for the problem. Plug in the indicated value of "x" as indicated by "f(x)".
.
For example: f(0) means to substitute "0" for "x" within the equation and solve.
a.) f(0)
f(x)=x^2-7x+10 [The frame-work]
f(0)=(0)^2-7(0)+10 [Substitute "0" for "x"]
f(0)=0-0+10 [Simplfy and solve]
f(0)=10 [Final answer]
.
.
f(x)=x2-7x+10
b.) f(5)
f(x)=x^2-7x+10 [The frame-work]
f(5)=(5)^2-7(5)+10 [Substitute "5" for "x"]
f(5)=25-35+10 [Simplfy and solve]
f(5)=0 [Final answer]
.
.
c.) f(1)
f(x)=x^2-7x+10 [The frame-work]
f(1)=(1)^2-7(1)+10 [Substitute "1" for "x"]
f(1)=1-7+10 [Simplfy and solve]
f(1)=4 [Final answer]